OpenStudy (anonymous):

An object is thrown up into the air on a windy day. The total flight time of the object is 10 seconds, and it travels 25 meters. Draw the trajectory of the rocket. How fast is the object traveling vertically upon takeoff??? Please help!

OpenStudy (anonymous):

It will travel 5 sec up and 5 sec down, so the height will be (1/2)(g)t^2 =(1/2)(10m/s^2)(25)=125 m. If it travels 25 m horizontally in 10 s, then its average velocity is 25/10=2.5 m/s and perhaps this is nearly constant. On the upward path, though, the velocity will be decreasing, v=vo-(10m/s^2)t and when it his 125m it will fall with increasing velocity, v=-(10m/s^2)(t-5). Trajectory should look like a parabola, a fly ball.